OpenStudy (anonymous):

6. The volume of a right circular cylinder is given by V =r2h. If a can has a radius of 3" and a height of 4", what volume of liquid will the can hold? V = pie symbal in3

OpenStudy (anonymous):

Just find the volume using formula

OpenStudy (anonymous):

if i knew how to do that i wouldnt be asking for help

OpenStudy (anonymous):

Okay, \[V=\pi r^2 h\] that's volume formula we know radius and height, lets plug them in \[\pi 3^2*4\] \[36\pi\]

OpenStudy (anonymous):

36 pi that's how much liquid it will hold

OpenStudy (anonymous):

V = in^3 this is what my answer box says so would it be 36 pie in^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you for ur help